Assume the required rate of return is 12 percent

Year                Cash Flows

0                      - $144,000

1                      $33,000

2                      $46,000

3                      $58,000

4                      $72,000

-What is the NPV of the above cash flows

- What is the IRR

-What is the Profitability Index

-What is the payback period of the project

-What is the discounted payback period
Answer #1

fax =SUM(D3:07) B 2 year Cash flows pv@12% 0$ (1,44,000.00) 4 1 $ 33,000.00 2 $ 46,000.00 3 $ 58,000.00 4 $ 72,000.00 IRR 14.

D8 for =SUM(D3:D7) pv@12% B 1 2 year Cash flows 30 -144000 4 1 33000 5 2 46000 63 58000 74 72000 8 IRR =IRR(B3:37) =C3/1.12 =